Brief interruption of authentication services today at 3:45PM

Please be advised that between 3:45PM and 4:00PM this afternoon, authentication services were briefly interrupted.  This resulted in several other services being unavailable during this period.  Services affected include Moodle, Campus Directory, VPN, Highlander Pipeline, etc.  The cause of this brief interruption is under investigation.  Please report any ongoing problems immediately to the IST ComputingContinue reading “Brief interruption of authentication services today at 3:45PM”

Status update at 6:30 pm (10/3/2010)

As of 5PM today all university-wide services,  with the exception of the Intellectual Property system (Inteum C/S),  are up and available.  The Intellectual Property system will be addressed tomorrow in conjunction with the Research & Development office.  All saved email was delivered by 6PM on Friday.  Please report abnormalities to the HelpDesk (973-596-2900, M-F 8am – 6pm). Thank youContinue reading “Status update at 6:30 pm (10/3/2010)”

Status update at 11:30 pm (10/2/2010)

All AFS data and services became available at 7pm today.  The following systems became available earlier today and were reported on previous postings:   Oak, HPC, Kong & Harp, AFS Student websites, database(s), AFS web services,  iTunes, Admissions Portal (applications), Student Employment (SEMS), Parking System (VRS), and Visitor Parking.

Status update at 4:00 pm (10/2/2010)

The following systems are now operational: iTunes, Admissions Portal (applications), Student Employment (SEMS), Parking System (VRS), and Visitor Parking. We continue to monitor the blog (https://njitsos.wordpress.com/) and answer as many of the questions as possible.   As further progress is made, that progress will be reported at this site. Thank you.

Status update at 1:40 pm (10/2/2010)

Work to restore all services continues 24 x 7.   Oak, HPC, Kong & Harp are available.  AFS Student websites, database(s)  & web services are available.  Work continues on the restoration of full services to Payroll Web Time Entry and Paystub Review, Admissions Portal and the remaining AFS services.  We continue to monitor the blog (https://njitsos.wordpress.com/) and answer as many of the questions asContinue reading “Status update at 1:40 pm (10/2/2010)”

Status update at 6:15 pm (10/1/2010)

 Numerous services were made available today.   These services include ADM email services for faculty and staff, Gmail authentication services for student email, DFS Network File Storage, Moodle courseware, many Highlander Pipeline applications (although several communication channels are not yet available), Alumni System, Budget Management System, Application Manager, Evisions, Online Catering, Visitor Parking, Eprint, etc.  Approximately 80% of AFS Network FileContinue reading “Status update at 6:15 pm (10/1/2010)”

Status update at 8:00 am

As of 5:00 am this morning, Oracle engineers have made the disk storage array available for NJIT applications and services.  At this time, ADM email and related services are available.  (ADM email (aka Outlook email) is typically used by NJIT faculty and staff.) Other services will be “re-started” and made available to the campus community. Continue reading “Status update at 8:00 am”

Status update at 8:00 pm

As of 6:00 pm this evening, Oracle engineers have successfully reconstructed all of the failed components of the disk storage array that has been out of service since early Tuesday morning. They are now in the process of reconnecting these components to make all of them available to the storage array’s controllers. This process willContinue reading “Status update at 8:00 pm”

Major service interruption update at 12:30 pm

As of noon today, the process to export 23 reconstructed volume groups was completed, a failed disk controller was replaced and now dual controllers in the storage array are functioning normally. Five of 18 volume groups needing to be reconstructed have been successfully reconstructed. Recovery procedures are run with due caution and continued success byContinue reading “Major service interruption update at 12:30 pm”

Major service interruption update 7:15 am

At midnight this morning, all automated data reconstruction was completed successfully on 23 of 41 volume groups comprising the disk array. The remaining 18 of 41 volume groups will be reconstructed in a less automated fashion, as was expected. A process to export the 23 reconstructed volume groups is nearly completed. This process serves toContinue reading “Major service interruption update 7:15 am”